1st- 40% passengers from a train got down at station A, 60% of the remaining passenger got down at station B, if there were still 540 passengers in the train, how many passengers were there before station A provided no one boarded the train at station A and station B?
2nd - 60% students in school are boys find the total number of students if 120 are girls?
3rd- A company reduced the price of an article by 5% by what percent the retailer must increase the new price so as to sell the article at the old price?
4th- anubhav bought 20 pens at rs 15 & 15 pens at rs 18 per pen . he sold all the pens at 17 rs per pen
find his gain or loss % in whole transaction?

Answer by mananth(16946)

before A -------------x
At A 40% got down
so balance 60% remained
=0.6x
At B 60$ of 0.6x got down
0.6 *0.6x= 0.36x
Balance = 0.6x-0.36x= 0.24x balance
0.24x=540
x=540/0.24
=2250 before A
----
let total be x
60% are boys
0.6x are boys
0.4x are girls
0.4x=120
x=120/0.4
x=300 total
.-------------
4th- anubhav bought 20 pens at rs 15 & 15 pens at rs 18 per pen . he sold all the pens at 17 rs per pen
find his gain or loss % in whole transaction?
20 pens Rs. 15 => 15*20 = 300
15 pens Rs. 18=> 15*18=270
Total cost = 300+270 = 570
...
Total pens 35 sold at Rs17
= 595
595-575= 20
he made Rs 20 profit on Rs 570 the cost price
so 100*20/570
= 3.5 % profit on cost